# Embodying Control in Soft Multistable Grippers from morphofunctional co-design

### v1

@@ A.3.2 Inverse problem $\rightarrow$ Position + Stiffness Optimization
A.3.3 Inverse problem $\rightarrow$ Position + Stiffness Optimization for multiple states
@@
$\min_{H_{1},H_{2},H_{3},H_{4},H_{5},H_{6},\text{Unit}_{L},\text{Unit}_{s}}w_{1}\sqrt{\left(\text{T1}_{xy}-\text{Tip dis 1}\right)^{2}}+w_{2}\sqrt{\left(\text{T2}_{xy}-\text{Tip dis 2}\right)^{2}}+w_{3}\sqrt{\left(\text{T3}_{xy}-\text{Tip dis 3}\right)^{2}}+w_{4}\frac{1}{\bar{\Pi}\left(R\right)}\$ (S.19)
Where $w_{1}=w_{2}=w_{3}=w_{4}=1/4$ and $T_{1}$ to $T_{3}$ represent the target positions for the gripper as shown in Table S3. Results for each of the targets can be seen in Fig. S7.
@@
A.3.4 Inverse problem $\rightarrow$ Bistable + Mestastable units
Our task planning application uses our dynamic model to calculate the resetting time of the metastable units. We fix our actuator geometry by selecting five segments with three bistable and two metastable units. The selected geometry parameters are $H_{1}=4$ mm, $H_{2}=4.5$ mm and $H_{3}=5$ mm, $\text{U}_{\text{L}}=9$ mm and $\text{U}_{\text{sep}}=1$ mm, and the last two unit heights ($H_{4}$ and $H_{5}$) is calculated to reset after 5 s with the pressure profile shown in Fig. S8a. Both dome heights are constrained to be the same ($H_{4}=H_{5}$), and the initial bistable set point can be observed in Fig. S8b. The full dynamic response predicted by the model is shown in Fig. S8c, where we can observe the three different snap-through instability and two snap-backs after 5s.
@@ A.3.4 Inverse problem $\rightarrow$ Bistable + Mestastable units
A.4 Experimental Validation
To validate our inverse design approach all five proposed geometries are 3D printed using Ninjaflex 85A. FDM printing is used given its versatility for printing complex geometries from flexible materials.
A.4.1 Inverse problem experimental validation
Five geometries are tested by activating all bistable unit cells and measuring the final tip position. The final position of the finger is measured by using image processing. Results can be observed in Fig. S9a (i-v), and the target coordinates, model prediction, and experimental measurements are reported in Table S4
Table S4: Experimental measurements for five different DPFs optimized for a specific target position.
@@ A.4.1 Inverse problem experimental validation
A.4.2 Experimental validation stiffness optimization
The stiffness of each DPF is measured using an Instron universal testing machine and applying a perpendicular force to the tip of the actuator (see Fig. S9b and Movie 4). Linear behavior is observed in the neighborhood of each stable state (0, 2, 4, and 5 active units in Fig. S9b), with a different stiffness for every state. Finally, the inverse design problem results are tested by comparing the stiffness of two geometries with the same target position but two distinct objective functions (Eq. 5 and Eq. 8).
Figure S9: Experimental results for different DPF geometries. a) Printed samples for inverse design target positions (i) - (v) b) Stiffness measurement for different active units. c) Comparison between DPF geometries for position inverse design and position+stiffness inverse design for target 5.

### v2

@@ A.6.2 Inverse problem $\rightarrow$ Position + Stiffness Optimization
t 0.83 0.86 0.81 0.85 0.83
$t_{\text{lim}}$ 1.5 1.36 1.45 1.35 1.4
UC 15.0 15.0 15.0 15.0 15.0
$K_{\text{st}}$ 0.011 0.0283 0.0146 0.0204 0.009
Improvement 1.22 1.13 1.46 1.13 1.3
Table S6: Inverse design problem results for five different targeted positions with maximum stiffness (Cheetah 95A).
Target 1 Target 2 Target 3 Target 4 Target 5
[46 32] [25 20] [25 40] [25 30] [35 45]
H1 4.99 4.66 4.96 5.0 4.9
H2 4.88 4.41 4.9 4.78 4.687
H3 4.86 4.31 4.72 4.69 4.57
H4 4.15 4.4 4.66 4.225
H5 3.92 4.2
H6
Unit S 2.26 1.4 1.11 1.01 1.03
Unit L 9.98 6.85 7.11 6.89 8.92
t 0.9 0.91 0.88 0.99 0.99
$t_{\text{lim}}$ 1.89 1.95 1.24 1.66 1.94
UC 15.0 15.0 15.0 15.0 15.0
$K_{\text{st}}$ 0.093 0.341 0.079 0.112 0.063
Improvement 1.51 1.61 1.48 2.7 2.37
A.6.3 Inverse problem $\rightarrow$ Position + Stiffness Optimization for multiple states
@@
$\begin{split}\min_{H_{i},t,\text{U}^{i}_{L},\text{U}^{i}_{\text{sep}}}&\,w_{1}\left(\text{T1}_{xy}-\text{Tip}_{\text{dis 1}}\right)^{2}+w_{2}\left(\text{T2}_{xy}-\text{Tip}_{\text{dis 2}}\right)^{2}\\ &+w_{3}\left(\text{T3}_{xy}-\text{Tip}_{\text{dis 3}}\right)^{2}+w_{4}\left(\frac{K_{\text{dis}}}{dF_{\text{in}}/dx}\right)^{2}\end{split}$ (S.21)
s.t. $\quad H_{i+1}\geq H_{i},\quad i=1,\dots,N-1,$
$\quad U^{i+1}_{\text{sep}}=U^{i}_{\text{sep}},\quad i=1,\dots,N,$
$\quad U^{i+1}_{\text{L}}=U^{i}_{\text{L}},\quad i=1,\dots,N.$
@@
Where $w_{1}=w_{2}=w_{3}=w_{4}=1/4$ and $T_{1}$ to $T_{3}$ represent the target positions for the gripper as shown in Table S7. Results for each of the targets can be seen in Figure S12.
A.6.4 Inverse problem $\rightarrow$ Bistable + Mestastable units
Our task planning application uses our dynamic model to calculate the resetting time of the metastable units. We fix our actuator geometry by selecting six segments with four bistable and two metastable units. The selected geometry parameters are $H_{1}=5$ mm, $H_{2}=5$ mm and $H_{3}=5$ mm, $\text{U}_{\text{L}}=6.5$ mm and $\text{U}_{\text{sep}}=1$ mm, $t=0.8$ mm, and the last two-unit heights ($H_{5}$ and $H_{6}$) is designed to maximize the resetting time under the two different pressure profiles show in Figure S13a. Both dome heights are constrained to be the same ($H_{5}=H_{6}$), and the initial bistable set point can be observed in Figure S13b. The full dynamic response predicted by the model is shown in Figure S13c, where we can observe the three different snap-through instability and two snap-backs in the two simulated cases. A restriction is imposed over these units to guarantee that they were metastable, by using the state map reported Figure S1b. As expected, to maximize the resetting time at two different conditions, the algorithm would converge to the highest possible metastable unit (see Figure S1b).
@@ A.6.4 Inverse problem $\rightarrow$ Bistable + Mestastable units
A.7 Experimental Test
A.7.1 Pressure and displacements Measurements
As shown in Figure S14, an air compressor (Pump DOA-P704-AA) is used to supply air to the DPF. The pressure is measured using a HONEYWELL ABPDANN010BG2A3 pressure sensor (0-10 Bar range), which is connected to an Arduino Uno for data acquisition. The Arduino records pressure readings at 500ms intervals to ensure accurate monitoring. The pressure is gradually increased to a specified value before being released to invert the dome units. The dynamic behavior of the DPF is captured in detail using a Photron Fastcam Mini UX100 high-speed camera, as illustrated in Movie 1. Tip displacement over time is then analyzed using the Tracker Video Analysis and Modeling Tool, providing precise measurements of the system’s response.
Figure S14: Experimental setup for pressure and displacement measurement. Schematic of the test setup used to record the displacement and pressure over time of the DPF and DFR. (1) DPR or DPF (2) Air Pump. (3) Pressure sensor. (4) Arduino Uno. (5) Laptop. (6) High Photron Speed camera.
A.7.2 Inverse problem experimental validation
Five geometries are tested by activating all bistable unit cells and measuring the final tip position. The final position of the finger is measured by using image processing. Results can be observed in Figure S15a (i-v) for NinjaFlex 85A and Figure S15b (i-iii) for Cheetah 95A, and the target coordinates, model prediction, and experimental measurements are reported in Table S9
Figure S15: Experimental results for different DPF geometries. a) Printed samples for inverse design target positions (i) - (v) for Ninjaflex 85A b) Printed samples for inverse design target positions (i) - (iii) for Cheetah 95A.
@@ A.7.2 Inverse problem experimental validation
Table S8: Experimental measurements for five different DPFs optimized for a specific target position (NinjaFlex 85A).
Target Coordinates Model Experiments Error
[46 , 32] [ 45.6, 31.8 ] [43.31 , 28.38] 6.9%
[25 , 20] [ 24.7, 19.5] [ 25,17, 16.55] 4.3%
[25 , 40] [ 25.15, 39.3] [ 23.03, 39.29] 2.5%
[25 , 30] [25.2, 29.22] [ 22.9, 31.6] 1.1%
[35 , 45] [ 33.97, 45.3] [ 33.35, 44.9] 1.2%
Table S9: Experimental measurements for five different DPFs optimized for a specific target position (Cheetah 95A).
A.7.3 Experimental validation stiffness optimization
The stiffness of each DPF is measured using an Instron universal testing machine and applying a perpendicular force to the tip of the actuator (see Figure S15b and Movie 5). Linear behavior is observed in the neighborhood of each stable state (0, 2, 4, and 5 active units in Figure S15b), with a different stiffness for every state. Finally, the inverse design problem results are tested by comparing the stiffness of two geometries with the same target position but two distinct objective functions (Equation 6 and Equation 8).
Figure S16: Force vs displacement plot for different DPF. a) Experimental setup. b) Effect of number of active domes on DPF stiffness. c) Comparison between DPF geometries for position inverse design and position+stiffness inverse design for target 5 (Cheetah 95A and NinjaFlex 85A).
A.7.4 Durability Test
The robustness of the DPF design is tested by performing 100 cycles over three weeks to show the geometry’s repeatability and effect. Results show that after three weeks and 100 cycles per week, the tip position stays within 5% of the target tip displacement.
Figure S17: Durability test and time decay response for a) NinjaTek NinjaFlex b) NinjaTek Cheetah (95A).
Notice that the characteristics of multistability reduce the potential effects of material variability by making geometry (i.e., morphology) the dominant driver of our soft robot’s response.
A.7.5 Carrying Capacity
Payload capacity and object grasping tests were conducted to evaluate the performance of our DPR, demonstrate its grasping capabilities, and determine the maximum carrying weight of the system. The designed global states enable adjustments in output force and position based on the number of inverted domes (see Movie 6). This adaptability allows the DPR to grasp objects of various sizes and topologies by removing the actuation pressure. Figure S18 illustrates the range of objects successfully grasped by the DPR using only the energy stored in the structure due to dome inversion. As depicted, diverse object types are effectively grasped, highlighting the system’s versatility. Additionally, to characterize the payload capacity, weights were incrementally added to the object shown in Figure S18 (i) until the structure could no longer grasp it. A maximum load of 738 g was recorded, resulting in a maximum load-to-weight ratio of 10.25 (See Movie 6).
Figure S18: DPF grasping capabilities are shown with six different objects. a) Blackberry Replica, b) Tennis Ball, c) Toy Dinosaur, d) Baking Soda pack, e) Mini Fruit Can. f) Plush Toy g) Cylindrical Cream. h) Rubik’s Cube. i) Coffee Cup.
A.7.6 Scalability
As the mechanical response of the DPF is derived from the dome geometry and its stability, the DPF and robot can be scaled to increase to large dimensions just by considering the energy stored during the dome inversion. Following the analysis by Seffen and Vidoli [58], we can derive different additional relations for scaling, given the energy required for inverting a single dome. This energy is proportional to $\frac{Et^{3}}{1-\nu^{2}}\left(\frac{H}{R}\right)^{2}$ where E, $\nu$, t, and R are the elastic modulus, Poisson’s ratio, thickness, height, and radius of the domes respectively.
We can utilize the dome shallowness $\frac{H}{R}$ and the curvature-to-thickness ratio $\frac{t}{R}$ to either scale up or down the DPF geometry. To test this, we 3D printed two different geometries with the same energetics but a 2x difference.
Figure S19: DPF scalability test. a) Two different DPF specimens scale using adimensional relationships. b) Stable states with all dome units active and a comparison with our lattice model.